OpenStudy (anonymous):

What is the antiderivative of sin2x/cosx?

geerky42 (geerky42):

HINT: \(\sin(2x) = 2\sin(x)\cos(x)\)

OpenStudy (anonymous):

ok, so then 2sinxcosx/cosx?

OpenStudy (anonymous):

set u=cosx

geerky42 (geerky42):

You now can cancel \(\cos(x)\) out.

OpenStudy (anonymous):

oh yeah..

OpenStudy (anonymous):

so then it's just the antiderivative of 2sinx?

geerky42 (geerky42):

Yeah

OpenStudy (anonymous):

so it's -2cosx

geerky42 (geerky42):

+C

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you

geerky42 (geerky42):

No problem